### v1

# A Computational Phase Function Approach for Obtaining $\alpha\alpha$ Wavefunctions

Abstract
In this work, the phase function method (PFM) is applied for the first time to explicitly construct scattering wavefunctions for the $\alpha\alpha$ system using interaction parameters optimized from earlier phase shift analysis. While previous investigations employing PFM and related approaches have primarily focused on the reproduction of scattering phase shifts or cross sections, the present study advances the method to directly reconstruct physically meaningful radial scattering wavefunctions for the $\ell=0$, 2, and 4 partial waves without solving the Schrödinger equation. The computed wavefunctions exhibit well-behaved near-origin and asymptotic characteristics and show excellent agreement with the established resonating-group method results of Hiura et al.. This agreement confirms the numerical robustness and accuracy of the proposed framework. The present results establish PFM as an efficient, unified, and computationally attractive tool for scattering wavefunction reconstruction in cluster–cluster systems, opening new possibilities for its application to more complex nuclear scattering problems.
1 Introduction
Inverse problems occupy a central position in modern theoretical physics, particularly in quantum scattering theory, where one seeks to infer interaction properties from experimentally accessible observables. As noted by Kabanikhin, systematic investigations of inverse problems began in the 1950s in disciplines such as electrodynamics, quantum scattering, geophysics, and acoustics, and later expanded rapidly with the availability of high-performance computing resources [1]. Today, inverse scattering techniques play an important role across a wide range of scientific domains, including nuclear physics, where they provide a powerful framework for reconstructing interaction potentials, phase shifts, and related quantities from scattering data.
In recent years, renewed interest in heavy-ion and cluster scattering problems has motivated the development of advanced numerical and symbolic–numerical methods to address both direct and inverse formulations of the scattering problem. Notable contributions include the symbolic–numerical algorithm proposed by Gusev et al. [2] for inverse optical-model analysis and the numerical investigations of direct and inverse scattering for various interaction models by Puzynina and Vo Trong Thach [3]. These studies underscore the continuing relevance of inverse approaches in extracting physically meaningful information from scattering observables.
Within quantum scattering theory, the wave function provides the most complete description of an interaction process, as it contains all information necessary to determine measurable quantities such as scattering amplitudes, phase shifts, and cross sections. Experimentally, however, wave functions are not directly observable; instead, one measures asymptotic quantities such as differential and total cross sections, which are encoded in the scattering amplitude. The reconstruction of wave functions and amplitudes from scattering data therefore constitutes a fundamental challenge in nuclear reaction theory. Conventional approaches typically rely on solving the time-independent Schrödinger equation for a given interaction potential and extracting scattering information from the asymptotic behavior of the resulting wave functions. These methods can become computationally demanding, especially in the presence of long-range Coulomb interactions.
An alternative and efficient framework is provided by the variable phase approach (VPA), also known as the phase function method (PFM), which reformulates the scattering problem in terms of first-order nonlinear differential equations for the radial phase shift and amplitude. Unlike traditional $S$-matrix or Jost-function methods [16, 17], PFM requires only the interaction potential as input and avoids explicit construction of the full scattering wave function at the outset. The method has been successfully applied to a variety of nuclear systems, including neutron–proton [22, 23], proton–proton [24], neutron–deuteron [25, 26, 27], and $\alpha$$\alpha$ scattering [28]. More recently, Laha and collaborators have demonstrated the versatility of PFM in studying nucleon–nucleon, nucleon–nucleus, and nucleus–nucleus interactions using different phenomenological potentials [18, 19, 20, 21].
The $\alpha$$\alpha$ system represents an especially suitable testing ground for the phase function framework. The $\alpha$ particle is characterized by zero spin and isospin, a large binding energy of 28.3 MeV, and a small root-mean-square radius of about 1.44 fm, allowing it to be treated as an elementary cluster. Consequently, the 8Be nucleus can be viewed as a weakly bound two-$\alpha$ system, and $\alpha$$\alpha$ scattering may be modeled effectively as a two-body problem governed by a local interaction potential. Since the pioneering experimental studies of Rutherford and Chadwick in 1927, a substantial body of experimental phase shift data has become available [4, 5, 6, 7, 8], enabling detailed theoretical investigations of the $\alpha$$\alpha$ interaction.
Historically, a variety of phenomenological potentials have been proposed to describe $\alpha$$\alpha$ scattering. Early studies by Häfner [9] and subsequent extensions by Nilson and collaborators [8] demonstrated the importance of angular-momentum dependence in reproducing experimental phase shifts. Later analysis using optical-model approaches and complex Woods–Saxon interactions revealed that no single $\ell$-independent potential can account for all partial waves over a wide energy range [6, 7]. Among simpler models, Buck et al. [12] employed a Gaussian nuclear interaction supplemented by a Coulomb term represented through an error function,
$V(r)=-V_{a}e^{-\alpha r^{2}}+z_{1}z_{2}\frac{e^{2}}{r}\,\mathrm{erf}(\beta r),$ (1)
and achieved good agreement with experimental phase shifts for several even partial waves. Ali and Bodmer [13] proposed a two-term Gaussian interaction capable of reproducing low-energy phase shifts for selected $\ell$ values, while more elaborate Woods–Saxon forms incorporating inelastic effects were later introduced for higher energies [7].
While most previous investigations have focused primarily on reproducing scattering phase shifts or cross sections, comparatively less attention has been devoted to the explicit construction and analysis of radial scattering amplitudes and wave functions. The phase function framework offers a natural means to address this gap, as it provides simultaneous access to the radial phase shift, amplitude, and wave function within a unified formalism. Motivated by the success of molecular-type interactions in describing effective forces between composite systems, the present work employs the Morse potential as a model for the short-range nuclear interaction between two $\alpha$ clusters, with the long-range Coulomb repulsion incorporated via an error-function representation.
The principal objective of this paper is to obtain a unified description of $\alpha$$\alpha$ scattering in terms of wave functions and scattering amplitudes within the phase function framework. By employing an inverse approach constrained by high-precision experimental phase shift data, the radial evolution of the phase shift, amplitude, and wave function is determined for relevant partial waves. This study aims to demonstrate that PFM, combined with physically motivated interaction models, provides deeper insight into the dynamics of the $\alpha$$\alpha$ system beyond conventional phase-shift analysis.
2 Methodology
2.1 Morse Interaction Potential
In the present work, the short-range nuclear interaction between two $\alpha$ particles is modeled using the Morse potential, a well-established phenomenological form originally introduced to describe diatomic molecular interactions and later adopted in nuclear scattering studies [29]. The Morse potential is expressed as
$V_{\text{Morse}}(r)=V_{0}\left[e^{-2(r-r_{m})/a_{m}}-2e^{-(r-r_{m})/a_{m}}\right],$ (2)
where $V_{0}$ denotes the depth of the potential, $r_{m}$ represents the equilibrium separation at which the attractive interaction is maximal, and $a_{m}$ controls the range and diffuseness of the interaction.
The Morse potential exhibits several features that make it particularly suitable for modeling effective interactions between composite nuclear systems. It naturally incorporates a strong short-range repulsion, an attractive pocket at intermediate distances, and an exponentially decaying tail at large separations, all of which are characteristic of realistic nuclear forces. In addition, the time-independent Schrödinger equation with a Morse interaction is exactly solvable, leading to closed-form expressions for bound-state energies and wave functions. For scattering problems, the Morse potential admits analytical solutions for specific partial waves, most notably the $^{1}S_{0}$ state, a property not shared by many commonly used phenomenological interactions such as the Hulthén, Malfliet–Tjon, or Manning-Rosen potentials.
Another important advantage of the Morse potential is its shape invariance, which results in comparatively simpler radial wave functions and facilitates both analytical insight and numerical implementation. These features, combined with its demonstrated success in reproducing low-energy nucleon–nucleon and cluster scattering observables, motivate its adoption as the reference nuclear interaction in the present phase function analysis.
2.2 Coulomb Interaction
In addition to the short-range nuclear force, the long-range Coulomb repulsion between the two positively charged $\alpha$ particles is explicitly included. To account for the finite spatial extent of the $\alpha$ particle, the Coulomb interaction is represented using an error-function-modified form:
$V_{C}(r)=\frac{4e^{2}}{r}\,\mathrm{erf}(\beta r),$ (3)
where the error function is defined as
$\mathrm{erf}(\beta r)=\frac{2}{\sqrt{\pi}}\int_{0}^{\beta r}e^{-x^{2}}\,dx.$ (4)
The parameter $\beta$ is related to the root-mean-square (RMS) radius of the $\alpha$ particle through
$\beta=\frac{\sqrt{3}}{2R_{\alpha}}.$ (5)
This form of the Coulomb interaction represents an improvement over the point-charge or uniformly charged sphere approximations by smoothly regularizing the interaction at short distances. In the present calculations, the RMS radius of the $\alpha$ particle is taken as $R_{\alpha}=1.44$ fm , though $\beta$ may also be treated as an adjustable parameter in phase shift optimization. The combined use of the Morse nuclear potential and the improved Coulomb interaction provides a physically consistent and numerically stable interaction model for the $\alpha\alpha$ scattering system.
2.3 Phase Function Method (PFM)
The scattering of a spinless particle with energy $E$ and orbital angular momentum $\ell$ in a central potential $V(r)$ is governed by the radial Schrödinger equation
$\frac{\hbar^{2}}{2\mu}\left[\frac{d^{2}}{dr^{2}}+\left(k^{2}-\frac{\ell(\ell+1)}{r^{2}}\right)\right]u_{\ell}(k,r)=V(r)\,u_{\ell}(k,r),$ (6)
where $u_{\ell}(k,r)$ denotes the reduced radial wave function, $\mu$ is the reduced mass of the system, and $k=\sqrt{E/(\hbar^{2}/2\mu)}$ is the wave number. For the $\alpha$$\alpha$ system, the constant $\hbar^{2}/2\mu$ takes the value $10.44217~\text{MeV\,fm}^{2}$. Under nonrelativistic kinematics, the center-of-mass energy is related to the laboratory energy through $E_{\text{CM}}=\tfrac{1}{2}E_{\ell ab}$.
Instead of solving the second-order differential equation (6) for the wave function, the phase function method (PFM), also known as the variable phase approach, reformulates the scattering problem in terms of a first-order nonlinear differential equation for the phase shift. This transformation relies on the mathematical equivalence between linear second-order differential equations and first-order Riccati equations, as originally developed by Calogero and Babikov.
Within this framework, the phase shift $\delta_{\ell}(k,r)$ evolves with the radial coordinate according to
$\delta_{\ell}^{\prime}(k,r)=-\frac{V(r)}{k(\hbar^{2}/2\mu)}\left[\cos\!\left(\delta_{\ell}(k,r)\right)\hat{j}_{\ell}(kr)-\sin\!\left(\delta_{\ell}(k,r)\right)\hat{\eta}_{\ell}(kr)\right]^{2},$ (7)
where the prime denotes differentiation with respect to $r$, and $\hat{j}_{\ell}(kr)$ and $\hat{\eta}_{\ell}(kr)$ are the Riccati–Bessel and Riccati–Neumann functions, respectively. These functions are related to the Riccati–Hankel function of the first kind through $\hat{h}_{\ell}(kr)=-\hat{\eta}_{\ell}(kr)+i\,\hat{j}_{\ell}(kr)$.
Equation (7) may also be expressed in integral form as
$\delta_{\ell}(k,r)=-\frac{1}{k(\hbar^{2}/2\mu)}\int_{0}^{r}V(r^{\prime})\left[\cos\!\left(\delta_{\ell}(k,r^{\prime})\right)\hat{j}_{\ell}(kr^{\prime})-\sin\!\left(\delta_{\ell}(k,r^{\prime})\right)\hat{\eta}_{\ell}(kr^{\prime})\right]^{2}dr^{\prime}.$ (8)
[…]
A major advantage of PFM is that the phase shifts are obtained directly from the interaction potential without requiring explicit knowledge of the wave function. Consequently, the numerical task is reduced to solving a single first-order nonlinear differential equation rather than a second-order boundary-value problem.
For the $s$-wave ($\ell=0$), the Riccati functions simplify to $\hat{j}_{0}(kr)=\sin(kr)$ and $\hat{\eta}_{0}(kr)=-\cos(kr)$, leading to a particularly compact phase equation,
$\delta_{0}^{\prime}(k,r)=-\frac{V(r)}{k(\hbar^{2}/2\mu)}\sin^{2}\!\left[kr+\delta_{0}(k,r)\right].$ (9)
For higher partial waves, the centrifugal contribution introduces additional structure through higher-order Riccati functions. The corresponding phase equations for $\ell=2$ ($d$-wave) and $\ell=4$ ($g$-wave) are obtained by substituting the appropriate angular-momentum-dependent forms of $\hat{j}_{\ell}(kr)$ and $\hat{\eta}_{\ell}(kr)$ into Eq. (7). Similar expressions can be constructed for even higher partial waves using recurrence relations for the Riccati functions.
The phase equations are numerically integrated from the origin, subject to the initial condition $\delta_{\ell}(0)=0$, up to sufficiently large distances where the interaction potential becomes negligible. In this asymptotic region, the phase function saturates to a constant value corresponding to the physical scattering phase shift. In the present work, the integration is carried out using a Runge–Kutta scheme of order four or five, ensuring numerical stability and accuracy across the full energy range considered.
Figure 1: Detailed flowchart to obtain scattering phase shift $\delta(E)$, amplitude $A(r)$ and wavefunction $u(r)$ for $\ell=0$, 2 & 4 waves.
2.3.1 Phase equations for $s$, $d$, and $g$ waves
Within the phase function framework, the scattering phase shift for each partial wave evolves according to a first-order nonlinear differential equation. By substituting the appropriate Riccati–Bessel and Riccati–Neumann functions into the general phase equation, explicit expressions are obtained for different angular momentum channels.
$s$-wave ($\ell=0$):
For the $s$-wave, the Riccati functions reduce to $\hat{j}_{0}(kr)=\sin(kr)$ and $\hat{\eta}_{0}(kr)=-\cos(kr)$. Consequently, the phase equation simplifies to
$\delta_{0}^{\prime}(k,r)=-\frac{V(r)}{k(\hbar^{2}/2\mu)}\sin^{2}\!\left[kr+\delta_{0}(k,r)\right]$ (10)
$d$-wave ($\ell=2$):
For the $d$-wave, the centrifugal term introduces higher-order radial contributions. The corresponding phase equation is given by
$\delta_{2}^{\prime}(k,r)=-\frac{V(r)}{k(\hbar^{2}/2\mu)}\left[-\sin\!\left(kr+\delta_{2}\right)-\frac{3}{kr}\cos\!\left(kr+\delta_{2}\right)+\frac{3}{(kr)^{2}}\sin\!\left(kr+\delta_{2}\right)\right]^{2}$ (11)
$g$-wave ($\ell=4$):
For the $g$-wave, the phase equation involves higher-order angular momentum terms and takes the form
$\delta_{4}^{\prime}(k,r)$ $=-\frac{V(r)}{k(\hbar^{2}/2\mu)}\Bigg[\sin\!\left(kr+\delta_{4}\right)+\frac{10}{kr}\cos\!\left(kr+\delta_{4}\right)-\frac{45}{(kr)^{2}}\sin\!\left(kr+\delta_{4}\right)$ (12)
$\hskip 73.97733pt-\frac{105}{(kr)^{3}}\cos\!\left(kr+\delta_{4}\right)+\frac{105}{(kr)^{4}}\sin\!\left(kr+\delta_{4}\right)\Bigg]^{2}.$
These equations are numerically integrated from the origin with the initial condition $\delta_{\ell}(0)=0$ up to the asymptotic region where the interaction potential vanishes. The saturation value of $\delta_{\ell}(k,r)$ at large $r$ yields the physical scattering phase shift for the corresponding partial wave.
2.4 Amplitude function A(r)
Within the phase function formalism, the radial amplitude $A_{\ell}(r)$ evolves with distance under the combined influence of the interaction potential and the running phase shift. The first-order differential equation governing the amplitude incorporates the local phase $\delta_{\ell}(k,r)$ and the Riccati–Bessel functions $\hat{j}_{\ell}(kr)$ and $\hat{\eta}_{\ell}(kr)$, thereby ensuring consistency with the partial-wave decomposition of the scattering solution. The general expression for the amplitude derivative is given by
$A_{\ell}^{\prime}(r)=$ $-\frac{A_{\ell}V(r)}{k}\left[\cos\left(\delta_{\ell}(k,r)\right)\hat{j}_{\ell}(kr)-\sin\left(\delta_{\ell}(k,r)\right)\hat{\eta}_{\ell}(kr)\right]$ (13)
$\times\left[\sin\left(\delta_{\ell}(k,r)\right)\hat{j}_{\ell}(kr)+\cos\left(\delta_{\ell}(k,r)\right)\hat{\eta}_{\ell}(kr)\right],$
Amplitude function equation for $\ell=0$ is given as
$A_{0}^{\prime}$ $=-\frac{A_{0}V(r)}{k\left(\frac{\hbar^{2}}{2\mu}\right)}\left[\cos\delta_{0}\sin(kr)+\sin\delta_{0}\cos(kr)\right]$
$\quad\times\left[\sin\delta_{0}\sin(kr)-\cos\delta_{0}\cos(kr)\right]$ (15)
Amplitude function equation for $\ell=2$ is given as
$A_{2}^{\prime}$ $=-\frac{A_{2}V(r)}{k\left(\frac{\hbar^{2}}{2\mu}\right)}\Bigg[\cos\delta_{2}\left(\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{(kr)}\cos(kr)\right)$
$\quad\times\Bigg(\sin\delta_{2}\left(\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{(kr)}\cos(kr)\right)$
$\qquad+\cos\delta_{2}\left(\left(-\frac{3}{(kr)^{2}}+1\right)\cos(kr)-\frac{3}{(kr)}\sin(kr)\right)\Bigg)\Bigg]$ (16)
Finally amplitude function equation for $\ell=4$ is given as
$A_{4}^{\prime}$ $=-\frac{A_{4}V(r)}{k\left(\frac{\hbar^{2}}{2\mu}\right)}\Bigg[\cos\delta_{4}\Bigg(\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\frac{10-105(kr)^{-2}}{kr}\cos(kr)\Bigg)$
$\quad-\sin\delta_{4}\Bigg(\left(-\frac{105}{(kr)^{2}}+10\right)\frac{\sin(kr)}{kr}+\left(-1+\frac{45}{(kr)^{2}}-\frac{105}{(kr)^{4}}\right)\cos(kr)\Bigg)\Bigg]$
$\quad\times\Bigg[\sin\delta_{4}\Bigg(\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\frac{10-105(kr)^{-2}}{kr}\cos(kr)\Bigg)$
$\qquad+\cos\delta_{4}\Bigg(\left(-\frac{105}{(kr)^{2}}+10\right)\frac{\sin(kr)}{kr}+\left(-1+\frac{45}{(kr)^{2}}-\frac{105}{(kr)^{4}}\right)\cos(kr)\Bigg)\Bigg]$ (17)
For the $s$-wave ($\ell=0$), the above equation simplifies due to the elementary forms of the spherical Bessel functions, leading to a comparatively compact expression for $A_{0}^{\prime}(r)$. The corresponding amplitude equation explicitly reflects the interplay between the potential strength and the running phase shift. A similar procedure is adopted for higher partial waves. In particular, the amplitude evolution for $\ell=2$ involves higher-order angular momentum terms arising from the centrifugal contribution, resulting in a more structured radial dependence. The amplitude equations for $\ell=0$ and $\ell=2$ are presented explicitly, while the corresponding expression for $\ell=4$ can be obtained analogously by employing the appropriate Riccati–Bessel functions.
The radial variation of the wavefunctions for $\ell=0$, $2$, and $4$ is illustrated in Fig. LABEL:fig2, highlighting the gradual stabilization of $A_{\ell}(r)$ at large distances where the interaction becomes negligible.
2.5 Wave function u(r)
Once the amplitude and phase functions are determined, the radial wave function for each partial wave can be constructed directly. Within the phase function framework, the reduced radial wave function is expressed as
$u_{\ell}(r)=A_{\ell}(r)\left[\cos\left(\delta_{\ell}(k,r)\right)\hat{j}_{\ell}(kr)-\sin\left(\delta_{\ell}(k,r)\right)\hat{\eta}_{\ell}(kr)\right],$ (18)
which ensures the correct asymptotic behavior of the scattering solution.
Wavefunction equation for $\ell=0$ is given as
$u_{0}(r)$ $=A_{0}(r)\left[\cos\delta_{0}(r)\,\sin(kr)-\sin\delta_{0}(r)\,\cos(kr)\right]$ (20)
Wavefunction equation for $\ell=2$ is given as
$u_{2}(r)$ $=A_{2}(r)\Bigg[\cos\delta_{2}(r)\left(\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{(kr)}\cos(kr)\right)$
$\quad-\sin\delta_{2}(r)\left(\left(-\frac{3}{(kr)^{2}}+1\right)\cos(kr)+\frac{3}{(kr)}\sin(kr)\right)\Bigg]$ (21)
Finally wavefunction equation for $\ell=4$ is given as
$u_{4}(r)$ $=A_{4}(r)\Bigg[\cos\delta_{4}(r)\Bigg(\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\frac{10-105(kr)^{-2}}{kr}\cos(kr)\Bigg)$
$\quad-\sin\delta_{4}(r)\Bigg(\left(-\frac{105}{(kr)^{2}}+10\right)\frac{\sin(kr)}{kr}+\left(-1+\frac{45}{(kr)^{2}}-\frac{105}{(kr)^{4}}\right)\cos(kr)\Bigg)\Bigg]$ (22)
For $\ell=0$, the wave function reduces to a simple combination of sine and cosine functions modulated by the amplitude $A_{0}(r)$ and the running phase shift $\delta_{0}(r)$. In the case of $\ell=2$, the centrifugal barrier introduces additional radial structure through higher-order trigonometric terms, resulting in a more oscillatory behavior of the wave function at short distances. Explicit expressions for the $s$- and $d$-wave radial wave functions are provided, while the formulation for $\ell=4$ follows straightforwardly by substituting the corresponding spherical Bessel functions.
The radial dependence of the wave functions for $\ell=0$, $2$, and $4$ is displayed in Fig. 2, demonstrating the smooth transition from interaction-dominated behavior at small $r$ to asymptotic free-particle motion at larger separations.
3 Results and Discussion
Using the optimized interaction parameters obtained from earlier phase-shift analysis, scattering wavefunctions for the $\ell=0$, 2, and 4 partial waves of the $\alpha\alpha$ system have been computed within the phase function framework. The calculations demonstrate that the phase function method enables a stable and consistent reconstruction of scattering wavefunctions without directly solving the second-order Schrödinger equation. The resulting radial wavefunctions exhibit smooth behavior throughout the interaction region and approach the correct asymptotic form at large distances, confirming that the underlying phase and amplitude functions are accurately determined. The gradual saturation of the amplitude function further indicates that the interaction effects are fully incorporated within the finite radial domain considered.
Figure 2: Wavefunction for s, p & d-waves at resonance energies in comparison with results of Hiura et. al.
For the s-wave ($\ell=0$), the wavefunction displays the expected monotonic behavior at short distances and a well-defined oscillatory structure at larger separations, reflecting the dominance of the nuclear interaction at low angular momentum. In the case of the d-wave ($\ell=2$), the presence of the centrifugal barrier introduces additional radial structure, leading to suppressed amplitudes at small radii and enhanced oscillatory behavior beyond the interaction region. The g-wave ($\ell=4$) exhibits an even stronger centrifugal effect, resulting in comparatively smaller amplitudes and a delayed onset of asymptotic oscillations. These trends are consistent with the expected angular-momentum dependence of cluster–cluster scattering and validate the numerical stability of the method across higher partial waves.
An important outcome of the present analysis is that the wavefunctions obtained using the phase function formalism retain physical transparency while remaining computationally efficient. Since the phase and amplitude functions are evolved simultaneously, the method provides direct access to the internal structure of the scattering solution, offering insight into how the interaction potential shapes the radial behavior of the wavefunction. The use of a Morse-type nuclear interaction combined with a finite-size Coulomb term ensures that both short-range and long-range effects are treated in a unified manner, leading to well-behaved solutions over the entire radial domain.
Overall, the results confirm that the phase function method is not limited to the determination of scattering phase shifts alone but can be reliably extended to construct scattering wavefunctions for the $\alpha\alpha$ system. The approach offers a compact and numerically stable framework for studying cluster scattering, making it particularly suitable for applications where computational efficiency and direct control over scattering observables are essential.
Figure 3: Wavefunction for s-wave at 19.45 MeV in comparison with results of Yoshiharu et. al.
4 Conclusion
In the present work, the phase function method has been successfully extended beyond its conventional application to scattering phase shifts in order to construct scattering wavefunctions for the $\alpha\alpha$ system. By employing optimized interaction parameters obtained from earlier phase-shift analyses, wavefunctions for the $\ell=0$, 2, and 4 partial waves have been computed within a unified and numerically stable framework. The formulation avoids the direct solution of the second-order Schrödinger equation and instead relies on first-order differential equations for the phase and amplitude functions, thereby reducing computational complexity while preserving the essential physical content of the scattering problem.
The resulting wavefunctions exhibit physically consistent behavior over the entire radial domain, including smooth evolution in the interaction region and correct asymptotic oscillatory behavior at large distances. The angular-momentum dependence of the solutions is clearly manifested through the increasing influence of the centrifugal barrier from s- to g-waves, demonstrating that the method reliably captures the characteristic features of cluster–cluster scattering dynamics. The stabilization of the amplitude function at large radii further confirms the internal consistency of the approach and the suitability of the adopted interaction model.
A significant outcome of the present analysis is that the computed wavefunctions show very good agreement with previously reported results by Hiura et al., thereby providing an important validation of the present phase-function-based formulation. This agreement indicates that the essential physics of the $\alpha\alpha$ interaction is effectively encoded in the optimized potential and the phase evolution, even without explicitly invoking more elaborate microscopic formalisms. The consistency with established results strengthens confidence in the accuracy and reliability of the present computational approach.
The use of a Morse-type nuclear interaction in conjunction with a finite-size Coulomb potential ensures a realistic and well-behaved description of both short-range nuclear attraction and long-range electrostatic repulsion. This combination contributes to the numerical stability of the solutions and allows for a transparent interpretation of the radial structure of the scattering wavefunctions. The ability to reconstruct wavefunctions directly from phase evolution highlights the strength of the phase function method as a unified tool for scattering studies.
Overall, this investigation demonstrates that the phase function framework provides an efficient and physically meaningful approach for analyzing $\alpha\alpha$ scattering, yielding scattering amplitudes and wavefunctions in addition to phase shifts within a single formalism. The methodology developed here can be readily extended to other cluster and nucleus–nucleus scattering systems and to different interaction models, making it a promising tool for future studies of nuclear scattering and inverse problems.
Data and Code Availibility
The numerical data and computational codes used in the study (including simulation and analysis routines) are available from the corresponding author upon reasonable request.

### v2

@@
-A Computational Phase Function Approach for Obtaining $\alpha\alpha$ Wavefunctions
+A Computational Phase Function Method for $\alpha-\alpha$ Scattering: Wavefunction Construction from Single and Two-Term Morse Potentials
 Abstract
-In this work, the phase function method (PFM) is applied for the first time to explicitly construct scattering wavefunctions for the $\alpha\alpha$ system using interaction parameters optimized from earlier phase shift analysis.
-While previous investigations employing PFM and related approaches have primarily focused on the reproduction of scattering phase shifts or cross sections, the present study advances the method to directly reconstruct physically meaningful radial scattering wavefunctions for the $\ell=0$, 2, and 4 partial waves without solving the Schrödinger equation.
-The computed wavefunctions exhibit well-behaved near-origin and asymptotic characteristics and show excellent agreement with the established resonating-group method results of Hiura et al..
-This agreement confirms the numerical robustness and accuracy of the proposed framework.
-The present results establish PFM as an efficient, unified, and computationally attractive tool for scattering wavefunction reconstruction in cluster–cluster systems, opening new possibilities for its application to more complex nuclear scattering problems.
+[…]
+For comparison, we adopt the interaction potential parameters reported by Sastri et al., who determined them using a two-term reference potential approach with genetic algorithm optimization to accurately reproduce the $\alpha\alpha$ scattering phase shifts.
+Without re-optimization, we construct the corresponding wavefunctions and find very good agreement with those obtained using our single-term Morse potential.
+The results also show excellent consistency with the resonating-group method calculations of Hiura et al..These findings demonstrate that PFM provides a numerically stable, efficient, and unified framework for scattering wavefunction construction in cluster–cluster systems.
 1 Introduction
@@
 These methods can become computationally demanding, especially in the presence of long-range Coulomb interactions.
-An alternative and efficient framework is provided by the variable phase approach (VPA), also known as the phase function method (PFM), which reformulates the scattering problem in terms of first-order nonlinear differential equations for the radial phase shift and amplitude.
-Unlike traditional $S$-matrix or Jost-function methods [16, 17], PFM requires only the interaction potential as input and avoids explicit construction of the full scattering wave function at the outset.
-The method has been successfully applied to a variety of nuclear systems, including neutron–proton [22, 23], proton–proton [24], neutron–deuteron [25, 26, 27], and $\alpha$$\alpha$ scattering [28].
-More recently, Laha and collaborators have demonstrated the versatility of PFM in studying nucleon–nucleon, nucleon–nucleus, and nucleus–nucleus interactions using different phenomenological potentials [18, 19, 20, 21].
-The $\alpha$$\alpha$ system represents an especially suitable testing ground for the phase function framework.
-The $\alpha$ particle is characterized by zero spin and isospin, a large binding energy of 28.3 MeV, and a small root-mean-square radius of about 1.44 fm, allowing it to be treated as an elementary cluster.
-Consequently, the 8Be nucleus can be viewed as a weakly bound two-$\alpha$ system, and $\alpha$$\alpha$ scattering may be modeled effectively as a two-body problem governed by a local interaction potential.
+[…]
+The $\alpha$$\alpha$ system represents an especially suitable testing ground for the phase function framework [22, 23].
+The $\alpha$ particle is characterized by zero spin and isospin, a large binding energy of 28.3 MeV, and a small root mean square radius of about 1.44 fm, allowing it to be treated as an elementary cluster.
+Consequently, the 8Be nucleus can be viewed as a weakly bound two $\alpha$ system, and $\alpha$-$\alpha$ scattering may be modeled effectively as a two-body problem governed by a local interaction potential.
 Since the pioneering experimental studies of Rutherford and Chadwick in 1927, a substantial body of experimental phase shift data has become available [24, 25, 26, 27, 28], enabling detailed theoretical investigations of the $\alpha$$\alpha$ interaction.
@@
 Early studies by Häfner [29] and subsequent extensions by Nilson and collaborators [28] demonstrated the importance of angular-momentum dependence in reproducing experimental phase shifts.
-Later analysis using optical-model approaches and complex Woods–Saxon interactions revealed that no single $\ell$-independent potential can account for all partial waves over a wide energy range [6, 7].
+Later analysis using optical-model approaches and complex Woods-Saxon interactions revealed that no single $\ell$-independent potential can account for all partial waves over a wide energy range [26, 27].
 Among simpler models, Buck et al.
@@
 By employing an inverse approach constrained by high-precision experimental phase shift data, the radial evolution of the phase shift, amplitude, and wave function is determined for relevant partial waves.
-This study aims to demonstrate that PFM, combined with physically motivated interaction models, provides deeper insight into the dynamics of the $\alpha$$\alpha$ system beyond conventional phase-shift analysis.
+This study aims to demonstrate that PFM [35, 36], combined with physically motivated interaction models, provides deeper insight into the dynamics of the $\alpha$$\alpha$ system beyond conventional phase-shift analysis.
 2 Methodology
@@
 In addition, the time-independent Schrödinger equation with a Morse interaction is exactly solvable, leading to closed-form expressions for bound-state energies and wave functions.
-For scattering problems, the Morse potential admits analytical solutions for specific partial waves, most notably the $^{1}S_{0}$ state, a property not shared by many commonly used phenomenological interactions such as the Hulthén, Malfliet–Tjon, or Manning-Rosen potentials.
+For scattering problems, the Morse potential admits analytical solutions for specific partial waves, most notably the $^{1}S_{0}$ state, a property not shared by many commonly used phenomenological interactions such as Manning-Rosen [19] potential, the Hulthén [21] potential or Malfliet-Tjon [38] potentials.
 Another important advantage of the Morse potential is its shape invariance, which results in comparatively simpler radial wave functions and facilitates both analytical insight and numerical implementation.
@@
 In addition to the short-range nuclear force, the long-range Coulomb repulsion between the two positively charged $\alpha$ particles is explicitly included.
-To account for the finite spatial extent of the $\alpha$ particle, the Coulomb interaction is represented using an error-function-modified form:
+To account for the finite spatial extent of the $\alpha$ particle, the Coulomb interaction is represented using an error-function-modified form [31]:
 $V_{C}(r)=\frac{4e^{2}}{r}\,\mathrm{erf}(\beta r),$ (3)
@@
 The combined use of the Morse nuclear potential and the improved Coulomb interaction provides a physically consistent and numerically stable interaction model for the $\alpha\alpha$ scattering system.
+2.3 Two-Component Morse Reference Potential
+The Reference Potential Approach (RPA), proposed by Selg [32, 33], provides a systematic procedure for treating one-dimensional quantum systems by selecting an analytically tractable reference interaction.
+Recently two-component Reference Potential Approach (RPA) was constructed from two smoothly connected Morse-type functions by Sastri et al.
+[34].
+The composite potential is defined in a piecewise form as
+$V_{\mathrm{RPA}}(r)=\begin{cases}V_{\mathrm{NC}}(r),&r\leq X,\\ V_{\mathrm{CL}}(r),&r\geq X,\end{cases}$ (6)
+where $X$ is the matching (boundary) radius.
+The inner-region potential is given by a standard Morse form,
+$V_{\mathrm{NC}}(r)=V_{1}+D_{1}\left[e^{-2\alpha_{1}(r-r_{1})}-2e^{-\alpha_{1}(r-r_{1})}\right],$ (7)
+and the outer-region potential is represented by a reversed Morse form,
+$V_{\mathrm{CL}}(r)=V_{2}-D_{2}\left[e^{-2\alpha_{2}(r-r_{2})}-2e^{-\alpha_{2}(r-r_{2})}\right].$ (8)
+Here,
+  • $D_{k}$ $(k=1,2)$ denote the depth parameters,
+  • $r_{k}$ are the equilibrium positions,
+  • $\alpha_{k}$ control the range (shape) of each Morse component,
+  • $V_{k}$ are constant offsets,
+  • $X$ is the matching radius.
+Continuity Conditions
+To ensure smooth matching at $r=X$, both the potential and its first derivative must be continuous:
+$V_{\mathrm{NC}}(X)=V_{\mathrm{CL}}(X),$ (9)
+$\left.\frac{dV_{\mathrm{NC}}}{dr}\right|_{X}=\left.\frac{dV_{\mathrm{CL}}}{dr}\right|_{X}.$ (10)
+Auxiliary Functions
+For compact representation, define
+$f_{1}$ $=e^{-2\alpha_{1}(X-r_{1})}-e^{-\alpha_{1}(X-r_{1})},$ (11)
+$f_{2}$ $=e^{-2\alpha_{2}(X-r_{2})}-e^{-\alpha_{2}(X-r_{2})},$ (12)
+$f_{3}$ $=e^{-2\alpha_{1}(X-r_{1})}-2e^{-\alpha_{1}(X-r_{1})},$ (13)
+$f_{4}$ $=e^{-2\alpha_{2}(X-r_{2})}-2e^{-\alpha_{2}(X-r_{2})}.$ (14)
+Determination of Depth Parameters
+Imposing the continuity conditions yields the following expressions for $D_{1}$ and $D_{2}$:
+$D_{1}=\frac{(V_{2}-V_{1})\alpha_{2}f_{2}}{\alpha_{2}f_{3}f_{2}-\alpha_{1}f_{4}f_{1}},$ (15)
+$D_{2}=\frac{(V_{1}-V_{2})\alpha_{1}f_{1}}{\alpha_{2}f_{3}f_{2}-\alpha_{1}f_{4}f_{1}}.$ (16)
+The resulting two-component Morse reference potential therefore depends on seven independent parameters:
+$V_{1},V_{2},\alpha_{1},\alpha_{2},r_{1},r_{2},X,$
+with $D_{1}$ and $D_{2}$ determined from the matching conditions.
+Sastri et al.
+determined the potential parameters using a genetic algorithm, which mitigates the risk of convergence to local minima.
+In the present work, we adopt the interaction potential constructed in Ref.
+[34] to compute the scattering phase shifts for the $\ell=0,2,$ and $4$ partial waves of $\alpha$$\alpha$ scattering.
+The results are systematically compared with those obtained from our single Morse potential formulation described in Sections 2.1 and 2.2.
 2.4 Phase Function Method (PFM)
@@
 Instead of solving the second-order differential equation (17) for the wave function, the phase function method (PFM), also known as the variable phase approach, reformulates the scattering problem in terms of a first-order nonlinear differential equation for the phase shift.
-This transformation relies on the mathematical equivalence between linear second-order differential equations and first-order Riccati equations, as originally developed by Calogero and Babikov.
+This transformation relies on the mathematical equivalence between linear second-order differential equations and first-order Riccati equations, as originally developed by Calogero [35] and Babikov [36].
 Within this framework, the phase shift $\delta_{\ell}(k,r)$ evolves with the radial coordinate according to
@@
 The saturation value of $\delta_{\ell}(k,r)$ at large $r$ yields the physical scattering phase shift for the corresponding partial wave.
-2.4 Amplitude function A(r)
-Within the phase function formalism, the radial amplitude $A_{\ell}(r)$ evolves with distance under the combined influence of the interaction potential and the running phase shift.
-The first-order differential equation governing the amplitude incorporates the local phase $\delta_{\ell}(k,r)$ and the Riccati–Bessel functions $\hat{j}_{\ell}(kr)$ and $\hat{\eta}_{\ell}(kr)$, thereby ensuring consistency with the partial-wave decomposition of the scattering solution.
-The general expression for the amplitude derivative is given by
-$A_{\ell}^{\prime}(r)=$ $-\frac{A_{\ell}V(r)}{k}\left[\cos\left(\delta_{\ell}(k,r)\right)\hat{j}_{\ell}(kr)-\sin\left(\delta_{\ell}(k,r)\right)\hat{\eta}_{\ell}(kr)\right]$ (13)
-$\times\left[\sin\left(\delta_{\ell}(k,r)\right)\hat{j}_{\ell}(kr)+\cos\left(\delta_{\ell}(k,r)\right)\hat{\eta}_{\ell}(kr)\right],$
-Amplitude function equation for $\ell=0$ is given as
-$A_{0}^{\prime}$ $=-\frac{A_{0}V(r)}{k\left(\frac{\hbar^{2}}{2\mu}\right)}\left[\cos\delta_{0}\sin(kr)+\sin\delta_{0}\cos(kr)\right]$
-$\quad\times\left[\sin\delta_{0}\sin(kr)-\cos\delta_{0}\cos(kr)\right]$ (15)
-Amplitude function equation for $\ell=2$ is given as
-$A_{2}^{\prime}$ $=-\frac{A_{2}V(r)}{k\left(\frac{\hbar^{2}}{2\mu}\right)}\Bigg[\cos\delta_{2}\left(\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{(kr)}\cos(kr)\right)$
-$\quad\times\Bigg(\sin\delta_{2}\left(\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{(kr)}\cos(kr)\right)$
-$\qquad+\cos\delta_{2}\left(\left(-\frac{3}{(kr)^{2}}+1\right)\cos(kr)-\frac{3}{(kr)}\sin(kr)\right)\Bigg)\Bigg]$ (16)
-Finally amplitude function equation for $\ell=4$ is given as
-$A_{4}^{\prime}$ $=-\frac{A_{4}V(r)}{k\left(\frac{\hbar^{2}}{2\mu}\right)}\Bigg[\cos\delta_{4}\Bigg(\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\frac{10-105(kr)^{-2}}{kr}\cos(kr)\Bigg)$
-$\quad-\sin\delta_{4}\Bigg(\left(-\frac{105}{(kr)^{2}}+10\right)\frac{\sin(kr)}{kr}+\left(-1+\frac{45}{(kr)^{2}}-\frac{105}{(kr)^{4}}\right)\cos(kr)\Bigg)\Bigg]$
-$\quad\times\Bigg[\sin\delta_{4}\Bigg(\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\frac{10-105(kr)^{-2}}{kr}\cos(kr)\Bigg)$
-$\qquad+\cos\delta_{4}\Bigg(\left(-\frac{105}{(kr)^{2}}+10\right)\frac{\sin(kr)}{kr}+\left(-1+\frac{45}{(kr)^{2}}-\frac{105}{(kr)^{4}}\right)\cos(kr)\Bigg)\Bigg]$ (17)
-For the $s$-wave ($\ell=0$), the above equation simplifies due to the elementary forms of the spherical Bessel functions, leading to a comparatively compact expression for $A_{0}^{\prime}(r)$.
-The corresponding amplitude equation explicitly reflects the interplay between the potential strength and the running phase shift.
-A similar procedure is adopted for higher partial waves.
-In particular, the amplitude evolution for $\ell=2$ involves higher-order angular momentum terms arising from the centrifugal contribution, resulting in a more structured radial dependence.
-The amplitude equations for $\ell=0$ and $\ell=2$ are presented explicitly, while the corresponding expression for $\ell=4$ can be obtained analogously by employing the appropriate Riccati–Bessel functions.
-The radial variation of the wavefunctions for $\ell=0$, $2$, and $4$ is illustrated in Fig.
-LABEL:fig2, highlighting the gradual stabilization of $A_{\ell}(r)$ at large distances where the interaction becomes negligible.
-2.5 Wave function u(r)
-Once the amplitude and phase functions are determined, the radial wave function for each partial wave can be constructed directly.
-Within the phase function framework, the reduced radial wave function is expressed as
-$u_{\ell}(r)=A_{\ell}(r)\left[\cos\left(\delta_{\ell}(k,r)\right)\hat{j}_{\ell}(kr)-\sin\left(\delta_{\ell}(k,r)\right)\hat{\eta}_{\ell}(kr)\right],$ (18)
-which ensures the correct asymptotic behavior of the scattering solution.
-Wavefunction equation for $\ell=0$ is given as
-$u_{0}(r)$ $=A_{0}(r)\left[\cos\delta_{0}(r)\,\sin(kr)-\sin\delta_{0}(r)\,\cos(kr)\right]$ (20)
-Wavefunction equation for $\ell=2$ is given as
-$u_{2}(r)$ $=A_{2}(r)\Bigg[\cos\delta_{2}(r)\left(\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{(kr)}\cos(kr)\right)$
-$\quad-\sin\delta_{2}(r)\left(\left(-\frac{3}{(kr)^{2}}+1\right)\cos(kr)+\frac{3}{(kr)}\sin(kr)\right)\Bigg]$ (21)
-Finally wavefunction equation for $\ell=4$ is given as
-$u_{4}(r)$ $=A_{4}(r)\Bigg[\cos\delta_{4}(r)\Bigg(\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\frac{10-105(kr)^{-2}}{kr}\cos(kr)\Bigg)$
-$\quad-\sin\delta_{4}(r)\Bigg(\left(-\frac{105}{(kr)^{2}}+10\right)\frac{\sin(kr)}{kr}+\left(-1+\frac{45}{(kr)^{2}}-\frac{105}{(kr)^{4}}\right)\cos(kr)\Bigg)\Bigg]$ (22)
-For $\ell=0$, the wave function reduces to a simple combination of sine and cosine functions modulated by the amplitude $A_{0}(r)$ and the running phase shift $\delta_{0}(r)$.
-In the case of $\ell=2$, the centrifugal barrier introduces additional radial structure through higher-order trigonometric terms, resulting in a more oscillatory behavior of the wave function at short distances.
-Explicit expressions for the $s$- and $d$-wave radial wave functions are provided, while the formulation for $\ell=4$ follows straightforwardly by substituting the corresponding spherical Bessel functions.
-The radial dependence of the wave functions for $\ell=0$, $2$, and $4$ is displayed in Fig.
-2, demonstrating the smooth transition from interaction-dominated behavior at small $r$ to asymptotic free-particle motion at larger separations.
+2.5 Amplitude Function $A_{\ell}(r)$
+Within the phase function framework, the scattering solution is expressed in terms of a running phase shift $\delta_{\ell}(k,r)$ and a radial amplitude function $A_{\ell}(r)$.
+While the phase equation governs the evolution of the local phase, the amplitude function determines the radial normalization of the scattering solution.
+Starting from the phase–amplitude decomposition of the reduced radial wave function,
+$u_{\ell}(r)=A_{\ell}(r)\Big[\cos\delta_{\ell}(k,r)\,\hat{j}_{\ell}(kr)-\sin\delta_{\ell}(k,r)\,\hat{\eta}_{\ell}(kr)\Big],$ (24)
+where $\hat{j}_{\ell}(kr)$ and $\hat{\eta}_{\ell}(kr)$ denote the Riccati–Bessel and Riccati–Neumann functions, respectively, one obtains the following first-order differential equation for the amplitude function:
+$A_{\ell}^{\prime}(r)=-\frac{V(r)}{k\left(\hbar^{2}/2\mu\right)}A_{\ell}(r)\Big[\cos\delta_{\ell}(k,r)\,\hat{j}_{\ell}(kr)-\sin\delta_{\ell}(k,r)\,\hat{\eta}_{\ell}(kr)\Big]\Big[\sin\delta_{\ell}(k,r)\,\hat{j}_{\ell}(kr)+\cos\delta_{\ell}(k,r)\,\hat{\eta}_{\ell}(kr)\Big].$ (25)
+This expression is valid for all partial waves. The amplitude equation is integrated simultaneously with the phase equation from the origin outward, subject to the initial condition
+$A_{\ell}(0)=1,$ (26)
+which ensures regular normalization at short distances.
+The evolution of $A_{\ell}(r)$ is driven by the interaction potential $V(r)$ and modulated by the local interference between regular and irregular free solutions.
+In the asymptotic region where $V(r)\rightarrow 0$, the derivative $A_{\ell}^{\prime}(r)\rightarrow 0$, and the amplitude approaches a constant value.
+This guarantees that the reconstructed wave function acquires the correct asymptotic scattering behavior.
+For the specific partial waves considered in the present work, the required Riccati functions are:
+S-wave ($\ell=0$)
+$\hat{j}_{0}(kr)=\sin(kr),\qquad\hat{\eta}_{0}(kr)=-\cos(kr).$ (27)
+D-wave ($\ell=2$)
+$\hat{j}_{2}(kr)=\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{kr}\cos(kr),$ (28)
+$\hat{\eta}_{2}(kr)=-\left(\frac{3}{(kr)^{2}}-1\right)\cos(kr)-\frac{3}{kr}\sin(kr).$ (29)
+G-wave ($\ell=4$)
+$\hat{j}_{4}(kr)=\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\left(\frac{10}{kr}-\frac{105}{(kr)^{3}}\right)\cos(kr),$ (30)
+$\hat{\eta}_{4}(kr)=-\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\cos(kr)+\left(\frac{10}{kr}-\frac{105}{(kr)^{3}}\right)\sin(kr).$ (31)
+Substitution of these expressions into the general amplitude equation yields the explicit forms employed in the numerical calculations.
+Retaining the compact Riccati representation preserves clarity and avoids unnecessary algebraic expansion.
+These are provided seperately in Appendix section.
+[…]
+2.6 Wave Function Construction
+Once the phase function $\delta_{\ell}(k,r)$ and the amplitude function $A_{\ell}(r)$ are determined, the reduced radial scattering wave function can be reconstructed directly within the phase function framework.
+The solution is expressed in the phase–amplitude form as
+$u_{\ell}(r)=A_{\ell}(r)\Big[\cos\delta_{\ell}(k,r)\,\hat{j}_{\ell}(kr)-\sin\delta_{\ell}(k,r)\,\hat{\eta}_{\ell}(kr)\Big],$ (32)
+where $\hat{j}_{\ell}(kr)$ and $\hat{\eta}_{\ell}(kr)$ denote the Riccati–Bessel and Riccati–Neumann functions, respectively.
+This representation ensures that the reconstructed wave function satisfies the correct asymptotic scattering boundary condition.
+In the asymptotic region $r\to\infty$, where the interaction potential $V(r)$ vanishes and the phase function saturates to its physical value $\delta_{\ell}(k)$, the wave function reduces to
+$u_{\ell}(r)\longrightarrow A_{\ell}(\infty)\sin\!\left(kr-\frac{\ell\pi}{2}+\delta_{\ell}(k)\right),$ (33)
+which is the standard form of the partial-wave scattering solution.
+For the specific partial waves considered in this work, the explicit expressions are obtained by substituting the corresponding Riccati functions.
+S-wave ($\ell=0$)
+$u_{0}(r)=A_{0}(r)\sin\!\big(kr+\delta_{0}(k,r)\big).$ (34)
+D-wave ($\ell=2$)
+$u_{2}(r)=A_{2}(r)\Big[\cos\delta_{2}(k,r)\,\hat{j}_{2}(kr)-\sin\delta_{2}(k,r)\,\hat{\eta}_{2}(kr)\Big].$ (35)
+G-wave ($\ell=4$)
+$u_{4}(r)=A_{4}(r)\Big[\cos\delta_{4}(k,r)\,\hat{j}_{4}(kr)-\sin\delta_{4}(k,r)\,\hat{\eta}_{4}(kr)\Big].$ (36)
+The radial behavior of $u_{\ell}(r)$ therefore reflects the combined evolution of the phase and amplitude functions.
+In the interaction region, deviations from free-particle oscillations encode the dynamical effects of the nuclear and Coulomb potentials, while at large distances the solution smoothly approaches the free asymptotic scattering form.
+The phase function framework thus provides a direct and numerically stable procedure for constructing scattering wave functions without explicitly solving the second-order Schrödinger equation.
+The method yields simultaneous access to phase shifts, amplitudes, and radial wave functions within a unified formalism.
 3 Results and Discussion
@@
 The gradual saturation of the amplitude function further indicates that the interaction effects are fully incorporated within the finite radial domain considered.
-Figure 2: Wavefunction for s, p & d-waves at resonance energies in comparison with results of Hiura et. al.
+Figure 2: (a) Scattering phase shifts and (b) interaction potentials for the $\ell=0$, 2, and 4 partial waves.
+The RPA results refer to the potentials constructed using the genetically optimized parameters reported by Sastri et al.
+[34].
+Figure 3: Variation of the phase shift (in radians) as a function of distance (in fm) for the $\ell=0$, 2, and 4 partial waves at resonant energy.
+The RPA results correspond to the reference potential approach using the parameters reported by Sastri et al.
+[34].
+Figure 4: Obtained scattering wavefunctions for S, P & D-waves in comparison with results of Hiura et.
+al.
+RPA refers to wavefunction obtained using reference potential approach parameters from Sastri et al.
+[34]
 For the s-wave ($\ell=0$), the wavefunction displays the expected monotonic behavior at short distances and a well-defined oscillatory structure at larger separations, reflecting the dominance of the nuclear interaction at low angular momentum.
@@
 The approach offers a compact and numerically stable framework for studying cluster scattering, making it particularly suitable for applications where computational efficiency and direct control over scattering observables are essential.
-Figure 3: Wavefunction for s-wave at 19.45 MeV in comparison with results of Yoshiharu et. al.
+Figure 5: Scattering wavefunction for S-wave at 19.45 MeV in comparison with results of Yoshiharu et al. [41]
 4 Conclusion
@@
 The resulting wavefunctions exhibit physically consistent behavior over the entire radial domain, including smooth evolution in the interaction region and correct asymptotic oscillatory behavior at large distances.
-The angular-momentum dependence of the solutions is clearly manifested through the increasing influence of the centrifugal barrier from s- to g-waves, demonstrating that the method reliably captures the characteristic features of cluster–cluster scattering dynamics.
+The angular-momentum dependence of the solutions is clearly manifested through the increasing influence of the centrifugal barrier from s- to g-waves, demonstrating that the method reliably captures the characteristic features of cluster-cluster scattering dynamics.
 The stabilization of the amplitude function at large radii further confirms the internal consistency of the approach and the suitability of the adopted interaction model.
-A significant outcome of the present analysis is that the computed wavefunctions show very good agreement with previously reported results by Hiura et al., thereby providing an important validation of the present phase-function-based formulation.
+A significant outcome of the present analysis is that the computed wavefunctions show very good agreement with previously reported results by Hiura et al., thereby providing an important validation of the present phase function based formulation.
 This agreement indicates that the essential physics of the $\alpha\alpha$ interaction is effectively encoded in the optimized potential and the phase evolution, even without explicitly invoking more elaborate microscopic formalisms.
@@
 The numerical data and computational codes used in the study (including simulation and analysis routines) are available from the corresponding author upon reasonable request.
+Appendix A Explicit Forms of the Amplitude and Wave Functions
+For completeness, we provide below the fully expanded expressions used in the numerical implementation for the $\ell=0,2,$ and $4$ partial waves.
+The equations follow directly from the general phase–amplitude formalism presented in Sec.
+2.
+A.1 Amplitude Equations
+The general amplitude equation is
+$A_{\ell}^{\prime}(r)=-\frac{A_{\ell}(r)\,V(r)}{k\left(\hbar^{2}/2\mu\right)}\Big[\cos\delta_{\ell}(r)\,\hat{j}_{\ell}(kr)-\sin\delta_{\ell}(r)\,\hat{\eta}_{\ell}(kr)\Big]\Big[\sin\delta_{\ell}(r)\,\hat{j}_{\ell}(kr)+\cos\delta_{\ell}(r)\,\hat{\eta}_{\ell}(kr)\Big].$ (37)
+Below we list the explicit forms.
+$\ell=0$
+Using
+$\hat{j}_{0}(kr)=\sin(kr),\qquad\hat{\eta}_{0}(kr)=-\cos(kr),$
+we obtain
+$A_{0}^{\prime}(r)$ $=-\frac{A_{0}V(r)}{k\left(\hbar^{2}/2\mu\right)}\left[\cos\delta_{0}\sin(kr)+\sin\delta_{0}\cos(kr)\right]$
+$\quad\times\left[\sin\delta_{0}\sin(kr)-\cos\delta_{0}\cos(kr)\right].$ (38)
+$\ell=2$
+For
+$\hat{j}_{2}(kr)$ $=\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{kr}\cos(kr),$
+$\hat{\eta}_{2}(kr)$ $=-\left(\frac{3}{(kr)^{2}}-1\right)\cos(kr)-\frac{3}{kr}\sin(kr),$
+the amplitude equation becomes
+$A_{2}^{\prime}(r)$ $=-\frac{A_{2}V(r)}{k\left(\hbar^{2}/2\mu\right)}\Big[\cos\delta_{2}\,\hat{j}_{2}(kr)-\sin\delta_{2}\,\hat{\eta}_{2}(kr)\Big]$
+$\quad\times\Big[\sin\delta_{2}\,\hat{j}_{2}(kr)+\cos\delta_{2}\,\hat{\eta}_{2}(kr)\Big].$ (39)
+$\ell=4$
+For
+$\hat{j}_{4}(kr)$ $=\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\left(\frac{10}{kr}-\frac{105}{(kr)^{3}}\right)\cos(kr),$
+$\hat{\eta}_{4}(kr)$ $=-\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\cos(kr)+\left(\frac{10}{kr}-\frac{105}{(kr)^{3}}\right)\sin(kr),$
+we obtain
+$A_{4}^{\prime}(r)$ $=-\frac{A_{4}V(r)}{k\left(\hbar^{2}/2\mu\right)}\Big[\cos\delta_{4}\,\hat{j}_{4}(kr)-\sin\delta_{4}\,\hat{\eta}_{4}(kr)\Big]$
+$\quad\times\Big[\sin\delta_{4}\,\hat{j}_{4}(kr)+\cos\delta_{4}\,\hat{\eta}_{4}(kr)\Big].$ (40)
+A.2 Wave Function Expressions
+The reconstructed reduced radial wave function is
+$u_{\ell}(r)=A_{\ell}(r)\Big[\cos\delta_{\ell}(r)\,\hat{j}_{\ell}(kr)-\sin\delta_{\ell}(r)\,\hat{\eta}_{\ell}(kr)\Big].$ (41)
+Explicitly:
+$\ell=0$
+$u_{0}(r)=A_{0}(r)\left[\cos\delta_{0}(r)\sin(kr)+\sin\delta_{0}(r)\cos(kr)\right].$ (42)
+$\ell=2$
+$u_{2}(r)=A_{2}(r)\Big[\cos\delta_{2}(r)\,\hat{j}_{2}(kr)-\sin\delta_{2}(r)\,\hat{\eta}_{2}(kr)\Big].$ (43)
+$\ell=4$
+$u_{4}(r)=A_{4}(r)\Big[\cos\delta_{4}(r)\,\hat{j}_{4}(kr)-\sin\delta_{4}(r)\,\hat{\eta}_{4}(kr)\Big].$ (44)